(y/x^2)+1+1/x dy/dx=0


1
Expert's answer
2021-06-28T16:48:40-0400

Let us solve the differential equation "\frac{y}{x^2}+1+\frac{1}{x} \frac{dy}{dx}=0," which is equivalent to "\frac{1}{x}y'+\frac{y}{x^2}=-1." Let us multiply both parts of the equation by "x^2:" "xy'+y=-x^2." The last equation is equivalent to "(xy)'=-x^2." It follows that "xy=-\frac{x^3}{3}+C." We conclude that the general solution is "y=-\frac{x^2}{3}+\frac{C}{x}."